6sqrt45 - 3sqrt80. I came up with 12sqrt3 as a answer and wasnt sure if that was right
Answer by dabanfield(803)

6sqrt45 - 3sqrt80. I came up with 12sqrt3 as a answer and wasnt sure if that was right
6*sqrt(45) - 3*sqrt(80) =
6*sqrt(9*5) - 3*sqrt(16*5) =
6*sqrt(9)*sqrt(5) - 3*sqrt(16)*sqrt(5) =
6*3*sqrt(5) - 3*4*sqrt(5) =
18*sqrt(5) - 12*sqrt(5) =
6*sqrt(5)